OpenStudy (anonymous):

URGENT PLEASE Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

b^2-4ac tells that if b^2-4ac>0, and is a perfect square then the roots can be found out by factoring in 2x^2=7x+3=0 we have 49-4*2*3=25>0 therefore roots can be found out by factorizing since 25 is aperfect sqre of 5 now 2x^2+7x+3=0 2x^2+6x+x+3=0 2x(x+3)+1(x+3)=0 u get (x+3)(2x+1)=0 x=-1/2,-3